A model is made of a car. The car is 9 feet long and the model is 6 inches long. What is the ratio of the length of the car to the length of the model?
A. 18:1
B. 1:18
C. 9:6
D. 6:9

Answer by user_dude2008(1825)

9 ft = 12*9 = 108 inches

108/6 = 18

Answer: A) 18:1